# Limit Laws for Poincaré Recurrence and the Shrinking Target Problem

### v1

@@ 1. Introduction
Following Boshernitzan’s work, Barreira and Saussol [1] and Urbánski [38] obtain a connection between recurrence rates and the pointwise dimension of the measure for various systems. More than 20 years since Boshernitzan’s pioneering work, several authors have improved the conclusion by imposing stricter conditions on the underlying systems, see for instance [30, 21, 23, 24].
@@
under some decay conditions on the sequence $M_{n}=\mu(B(x,r_{n}(x)))$. Persson and the author [34] also establish an error term. Considering radii $r_{n}$ that do not depend on the centre of the ball, He [15], when $\mu$ is absolutely continuous, and Levesley, Li, Simmons, Velani [27], when $\mu=m$ is the Lebesgue measure, obtain RSBCs for more general sets with error terms, while also relaxing conditions on the decay of $r_{n}$. In particular, we bring attention to the main result of He, which restricted to balls implies that
@@
One may view a Strong Borel–Cantelli lemma as a non-autonomous Strong Law of Large Numbers. As such, it is natural to ask if the sum in (1.2) satisfies a distributional law when normalized appropriately. Answering this question is the primary purpose of this paper. With this in mind, Theorem 3.3 in Conze and Le Borgne [9] can be applied to obtain a distributional law for the sum in (1.2) for a class of systems when $X$ is a compact group and $r_{k}(x)=r$ is constant in $x\in X$ and $k\geq 0$. More precisely, when $X$ is a compact group and the system $(X,T,\mu)$ satisfies a multiple decorrelation property, one can apply Conze and Le Borgne’s result to obtain that $x\mapsto n^{-1/2}\sum_{k=1}^{n}(\mathbf{1}_{B(x,r)}\circ T^{k}x-\mu(B(x,r)))$ converges in distribution to a random variable with characteristic function
$t\mapsto\int_{X}\exp\Bigl(-\frac{t^{2}}{2\sigma_{x}^{2}}\Bigr)\mathop{}\!d\mu(x),$
where $\sigma_{x}^{2}=\lim_{n\to\infty}n^{-1}\operatorname{Var}(\sum_{k=1}^{n}\mathbf{1}_{B(x,r)}\circ T^{k})$ is the asymptotic variance of $\mathbf{1}_{B(x,r)}$. This suggests that, in the more general setting, the sum in (1.2) satisfies a similar distributional law.
@@
Alternatively, Lévy’s continuity theorem ([22, Theorem 5.3]) implies that the condition (CLT) is equivalent to
@@
A stronger result than the CLT, is the almost sure invariance principle (ASIP), which can be stated simply as an almost sure version of the CLT. The particular version of the ASIP we are interested in is the following. We say that a sequence of random variables $(X_{n})$ satisfies the ASIP with rate $\beta\in(0,1)$ if there exists a sequence of standard Gaussian random variables $(Z_{n})$ (by enlarging the underlying probability space if necessary) such that
@@
and $\sum_{k=1}^{n}\mathbf{E}[Z_{k}^{2}]\sim\sigma_{n}^{2}$, where $\sim$ denotes asymptotic equivalence. The ASIP is a strong asymptotic estimate and implies other limit laws, such as a functional CLT and the law of the iterated logarithm (LIL).
@@
Theorems 2.3 and 2.5 extend that of Haydn, Nicol, Török and Vaienti [13, Theorem 6.1, Corollary 6.2] who employ an abstract framework due to Cuny and Merlevède [10, Theorem 2.3] (see Theorem 4.1), in order to prove (ASIP) for non-stationary Hölder continuous observables on non-invertible and invertible hyperbolic systems, respectively. The main drawback of their results is the requirement that the observables have uniformly bounded Hölder norm, which prevents the application of their results to the shrinking target problem. Nevertheless, the proofs of Theorems 2.3 and 2.5 are modifications of [13, Theorem 6.1, Corollary 6.2]. By improving their results, we also manage to prove a CLT for the shrinking target problem in the setting of uniformly hyperbolic systems, which hitherto had not been proven. We also note that Chen, Yang and Zhang [5] establish (ASIP) for non-uniformly hyperbolic systems equipped with the absolutely continuous invariant measure. We delay a more in-detail discussion of these results for when we state our results.
@@ 1.1. Structure of the Paper
We begin with establishing relevant notation and stating our assumptions in Section 2. We then state our results on the distributional law for the recurrence problem and the ASIP for non-autonomous observables, which is then applied to the shrinking target problem. In Section 3, we state various results on the behaviour of the variances that appear in the both the recurrence problem and shrinking target problem. Section 4 is devoted to stating various supporting results that are mostly found elsewhere. We devote Sections 58 to the proofs of the CLT, ASIP and variance results. Section 9 is devoted to a discussion on assumptions used in Theorem 2.1, Corollary 2.1 and variance Lemmas 3.4 and 3.6. More precisely, the section is devoted to establishing a short returns estimate when $\mu$ is absolutely continuous.
@@ 2.1. Assumptions
Throughout the paper, we assume that $(X,d)$ is a $\dim X$-dimensional smooth compact Riemannian manifold, and that the quintuplet $(X,T,\mu,\mathcal{F},d)$ is a metric measure-preserving dynamical system (m.m.p.s.). That is, $(X,\mu,\mathcal{F})$ is a probability space, $\mu$ is a Borel measure and $T\colon X\to X$ is a measurable transformation which preserves $\mu$. Unless relevant, we choose to suppress the mention of the $\sigma$-algebra $\mathcal{F}$. We denote the normalised transfer operator corresponding to the system by $P$, i.e., the $P\colon L^{2}(\mu)\to L^{2}(\mu)$ is the operator defined by $\int_{X}\varphi(\psi\circ T)\mathop{}\!d\mu=\int_{X}(P\varphi)\psi\mathop{}\!d\mu$ for all $\varphi,\psi\in L^{2}(\mu)$.
@@
Throughout the article, $(r_{n})$ will denote a sequence of radius functions $r_{n}\colon X\to(0,\infty)$ and we will consider the corresponding recurrence sets $E_{n}=\{x\in X:T^{n}x\in B(x,r_{n}(x))\}$, where $B(x,r)$ is the open ball with centre $x$ and radius $r$. We will predominantly consider two cases: one in which the functions $(r_{n})$ are defined implicitly with reference to a given sequence $(M_{n})$; and another in which $(r_{n})$ is a sequence of scalars, given without a reference to another sequence. More precisely, in the first case, given a sequence of real numbers $(M_{n})$ in $(0,1]$, define – whenever possible – the functions $(r_{n})$ by $\mu(B(x,r_{n}(x)))=M_{n}$ for all $x\in X$ and $n\geq 0$. We will broadly refer to the first case as the implicit radii problem, and to the second case as the explicit radii problem. In the latter problem, we will denote the recurrence sets as $\hat{E}_{n}=\{x\in X:T^{n}x\in B(x,r_{n})\}$ to clearly delineate between the problems. We will also denote related quantities with a hat. When considering explicit radii, we will find it useful to consider the sequence $(\hat{M}_{n})$ given by $\hat{M}_{n}=\int\mu(B(x,r_{n}))\mathop{}\!d\mu(x)$. In the implicit radii setting, we will see that the functions $r_{n}$ are well defined under certain assumptions on $\mu$ and that $r_{n}\to 0$ uniformly on $\operatorname{supp}\mu$. Note that, since $X$ is compact, we have that $r_{k}(x)\leq\lvert X\rvert$, whenever $r_{k}(x)$ is defined.
@@
The conditions on the measure imply that, when defined implicitly, $r_{n}$ is well defined for each $n\in\mathbb{N}$ and that $r_{n}$ Lipschitz continuous on $\operatorname{supp}\mu$ with $\lvert r_{n}\rvert_{\textup{Lip}}=1$. In the case that $M_{n}\to 0$, then $r_{n}\to 0$ uniformly on $\operatorname{supp}\mu$. (The proofs of these facts are straightforward [35, Lemmas 3.1, 3.2].) When $r_{n}\to 0$ uniformly on $\operatorname{supp}\mu$, we may therefore without loss of generality assume that $r_{n}\leq\rho_{0}$ $\mu$-a.e. for all $n\in\mathbb{N}$.
@@
Conditions M.1 and M.2 are standard assumptions; they are assumed, either implicitly or explicitly, in the aforementioned works establishing RSBCs [15, 20, 27, 33, 34, 35]. If $\mu$ is absolutely continuous with density $h\in L^{q}(m)$ for some $q>1$, then both conditions hold. Furthermore, if $X$ is a Riemannian manifold with dimension an $\dim X$ and $\mu$ satisfies Condition M.1 for $s>\dim X-1$, then Condition M.2 holds for $\alpha_{0}=\dim X+1-s$ (see the proof of Theorem 1.1 in [35]).
@@ 2.2. Distributional Limit Laws for Recurrence
Remark 4.
The condition in (2.3) that the ratio of variances converge in distribution can be verified if the system satisfies a short returns estimate of the following form: there exist $\delta_{0}>0$ and $\upsilon_{0}>1$ such that
@@
This will be proven in Lemma 3.4. The short returns estimate will be discussed further in Section 9.
Theorem 2.1 can be applied to interval maps equipped with a Gibbs measure, and Axiom A systems equipped with SRB measures satisfying (M.1) for $s>\dim X-1$.
We also establish a distributional limit in for the explicit radii problem. Recall that $\hat{E}_{n}=\{x\in X:T^{n}x\in B(x,r_{n})\}$ and $\hat{M}_{n}=\int\mu(B(x,r_{n}))\mathop{}\!d\mu(x)$. Let $\hat{\sigma}_{n}^{2}=\operatorname{Var}(\sum_{k=1}^{n}\mathbf{1}_{\hat{E}_{k}})$ and $\hat{s}_{n}^{2}=\operatorname{Var}(\sum_{k=1}^{n}\mathbf{1}_{B(y,r_{k})}\circ T^{k})$.
@@
Suppose that $(X,T,\mu,d)$ is an m.m.p.s. satisfying Condition D.3 and that $\mathop{}\!d\mu=h\mathop{}\!dm$ with density $h\in L^{2}(m)$ such that $h\geq c>0$ for some constant $c>0$. Assume that
@@
for all sufficiently large $k$ and all $l=1,\ldots,(\log k)^{p_{0}}$. If $(r_{n})$ satisfies Condition $\mathcal{S}_{\gamma,\upsilon}$ for some $\gamma\in(0,d)$ and $\upsilon>2(\delta_{0}\dim X)^{-1}$, then the characteristic function of
@@
Although the result holds for $h\in L^{2}(m)$ given that (2.6) holds, we can only verify (2.6) when $h\in L^{\infty}(m)$. See §9 for further details. Since $h$ is bounded away from 0, the limiting distribution has probability density function $f\colon\mathbb{R}\to(0,\infty)$ given by
$f(t)=\frac{1}{\sqrt{2\pi}}\int_{X}\frac{1}{\sqrt{h(y)}}\exp\Bigl(-\frac{t^{2}}{2h(y)}\Bigr)\mathop{}\!d\mu(y)$
@@
Remark 7.
Comparing with the result of Holland and Todd [20], namely (1.4), we see that both our probability density functions involve averaging over the ‘expected’ distributions. More precisely, the density in (1.4) is an average of Poisson distributions with variance $\tau h(x)$. In our case, our density is an average of Gaussian distributions with variance $h(x)$.
Comparing with the result of Conze and Le Borgne [9, Theorem 3.3], we recover their result when $X$ is a compact group and $r_{k}=r\in(0,\infty)$ for all $k$.
@@ 2.3. Almost Sure Invariance Principles for Non-invertible Systems
As mentioned in the introduction, Haydn, Nicol, Vaienti and Zhang [13] prove an ASIP (and therefore a CLT) for the shrinking target problem when the transfer operator of the system contracts in the $\mathrm{BV}$ norm, i.e., under Condition D.1 with $\|\cdot\|_{\alpha}$ replaced with $\|\cdot\|_{\textup{BV}}$. Thus, their result can be used to establish the conclusion of Theorem 2.1 for such systems; however, in order to prove the conclusion for a larger class of systems – including uniformly hyperbolic systems – one needs to weaken their assumptions. In [13, Theorem 6.1], the authors assume, for their sequence of observables $(\varphi_{n})$, that $\sup_{n\geq 1}\|\varphi_{n}\|_{\alpha}<\infty$. Therefore, they are unable to apply their result to the shrinking target problem. Theorem 2.3 relaxes this condition, which enables us to use an approximation argument and establish an ASIP for the shrinking target problem. Additionally, we improve [13, Theorem 6.1] by weakening their condition that $\sigma_{n}^{2}\geq n^{\delta}$ for some $\delta>\frac{\sqrt{17}-1}{4}\approx 0.781$ to $\sigma_{n}^{2}\geq n^{\delta}$ for any $\delta>0$. (We will see that $\sigma_{n}^{2}\gtrsim n^{1-\gamma}$.)

### v2

@@ 1. Introduction
Following Boshernitzan’s work, Barreira and Saussol [1] and Urbánski [42] obtain a connection between recurrence rates and the pointwise dimension of the measure for various systems. More than 20 years since Boshernitzan’s pioneering work, several authors have improved the conclusion by imposing stricter conditions on the underlying systems, see for instance [32, 23, 25, 26, 17].
@@
under some decay conditions on the sequence $M_{n}=\mu(B(x,r_{n}(x)))$. Persson and the author [36] also establish an error term. Considering radii $r_{n}$ that do not depend on the centre of the ball, He [16], when $\mu$ is absolutely continuous, and Levesley, Li, Simmons and Velani [29], when $\mu=m$ is the Lebesgue measure, obtain RSBCs for more general sets with error terms, while also relaxing conditions on the decay of $r_{n}$. In particular, we bring attention to the main result of He, which restricted to balls implies that
@@
One may view a Strong Borel–Cantelli lemma as a non-autonomous Strong Law of Large Numbers. As such, it is natural to ask if the sum in (1.2) satisfies a distributional law when normalized appropriately. Answering this question is the primary purpose of this paper. With this in mind, Theorem 3.3 in Conze and Le Borgne [9] states that if $X$ is a compact group, $\varphi$ is a Hölder continuous function, and $(X,T,\mu)$ satisfies a multiple decorrelation property (see Condition D.3 below), then $x\mapsto n^{-1/2}\sum_{k=1}^{n}\varphi(T^{k}x-x)-\mu(\varphi(\cdot-x))$ converges in distribution to a random variable with characteristic function
$t\mapsto\int_{X}\exp\Bigl(-\frac{1}{2}\sigma_{x}^{2}t^{2}\Bigr)\mathop{}\!d\mu(x),$
where $\sigma_{x}^{2}=\lim_{n\to\infty}n^{-1}\operatorname{Var}(\sum_{k=1}^{n}\varphi(T^{k}\cdot-x)-\mu(\varphi(\cdot-x)))$ is the asymptotic variance of $\varphi(\cdot-x)$. Taking $\varphi\approx\mathbf{1}_{B(0,r)}$ suggests that a similar result should hold for (1.2), at least when $r_{k}(x)=r$ for all $k\geq 1$ and $x\in X$.
@@
Alternatively, Lévy’s continuity theorem ([24, Theorem 5.3]) implies that the condition (CLT) is equivalent to
@@
A stronger result than the CLT is the almost sure invariance principle (ASIP), which can be stated simply as an almost sure version of the CLT. The particular version of the ASIP we are interested in is the following. We say that a sequence of random variables $(X_{n})$ satisfies the ASIP with rate $\beta\in(0,1)$ if there exists a sequence of standard Gaussian random variables $(Z_{n})$ (by enlarging the underlying probability space if necessary) such that
@@
and $\sum_{k=1}^{n}\mathbf{E}[Z_{k}^{2}]\sim\sigma_{n}^{2}$, where $\sim$ denotes asymptotic equivalence. The ASIP is a strong asymptotic estimate and implies other limit laws, such as a functional CLT and the law of the iterated logarithm (LIL) (see [37]).
@@
Theorems 2.3 and 2.5 extend that of Haydn, Nicol, Török and Vaienti [14, Theorem 6.1, Corollary 6.2] who employ an abstract framework due to Cuny and Merlevède [10, Theorem 2.3] (see Theorem 4.1), in order to prove (ASIP) for non-stationary Hölder continuous observables on non-invertible and invertible hyperbolic systems, respectively. The main drawback of their results is the requirement that the observables have uniformly bounded Hölder norm, which prevents the application of their results to the shrinking target problem. Nevertheless, the proofs of Theorems 2.3 and 2.5 are modifications of [14, Theorem 6.1, Corollary 6.2]. By improving their results, we also manage to prove a CLT for the shrinking target problem in the setting of uniformly hyperbolic systems, which hitherto had not been proven. We also note that Chen, Yang and Zhang [5] establish (ASIP) for non-uniformly hyperbolic systems equipped with the absolutely continuous invariant measure. We delay a more in-detail discussion of these results for when we state our results.
@@ 1.1. Structure of the Paper
We begin with establishing relevant notation and stating our assumptions in Section 2. We then state our results on the distributional law for the recurrence problem and the ASIP for non-autonomous observables, which is then applied to the shrinking target problem. In Section 3, we state various results on the behaviour of the variances that appear in the both the recurrence problem and shrinking target problem. Section 4 is devoted to stating various supporting results that are mostly found elsewhere. We devote Sections 58 to the proofs of the CLT, ASIP and variance results. Section 9 is devoted to a discussion on assumptions used in Theorem 2.1, Corollary 2.1 and variance Lemmas 3.4 and 3.6. More precisely, the section is devoted to establishing a short returns estimate when $\mu$ is absolutely continuous. We conclude with final remarks in Section 10, in which we mention further directions of research.
@@ 2.1. Assumptions
Throughout the paper, we assume that $(X,d)$ is a $(\dim X)$-dimensional smooth compact Riemannian manifold, and that the quintuplet $(X,T,\mu,\mathcal{F},d)$ is a metric measure-preserving dynamical system (m.m.p.s.). That is, $(X,\mu,\mathcal{F})$ is a probability space, $\mu$ is a Borel measure and $T\colon X\to X$ is a measurable transformation which preserves $\mu$. Unless relevant, we choose to suppress the mention of the $\sigma$-algebra $\mathcal{F}$. We denote the normalised transfer operator corresponding to the system by $P$, i.e., $P\colon L^{2}(\mu)\to L^{2}(\mu)$ is the operator defined by $\int_{X}\varphi(\psi\circ T)\mathop{}\!d\mu=\int_{X}(P\varphi)\psi\mathop{}\!d\mu$ for all $\varphi,\psi\in L^{2}(\mu)$.
@@
Throughout the article, $(r_{n})$ will denote a sequence of radius functions $r_{n}\colon X\to(0,\infty)$ and we will consider the corresponding recurrence sets $E_{n}=\{x\in X:T^{n}x\in B(x,r_{n}(x))\}$, where $B(x,r)$ is the open ball with centre $x$ and radius $r$. We will predominantly consider two cases: one in which the functions $(r_{n})$ are defined implicitly with reference to a given sequence $(M_{n})$; and another in which $(r_{n})$ is a sequence of scalars, given without a reference to another sequence. More precisely, in the first case, given a sequence of real numbers $(M_{n})$ in $(0,1]$, define – whenever possible – the functions $(r_{n})$ by $\mu(B(x,r_{n}(x)))=M_{n}$ for all $x\in X$ and $n\geq 0$. We will broadly refer to the first case as the implicit radii problem, and to the second case as the explicit radii problem. In the latter problem, we will denote the recurrence sets as $\hat{E}_{n}=\{x\in X:T^{n}x\in B(x,r_{n})\}$ to clearly delineate between the problems. We will also denote related quantities with a hat. When considering explicit radii, we will find it useful to consider the sequence $(\hat{M}_{n})$ given by $\hat{M}_{n}=\int\mu(B(x,r_{n}))\mathop{}\!d\mu(x)$. In the implicit radii setting, we will see that the functions $r_{n}$ are well defined under certain assumptions on $\mu$, and that $r_{n}\to 0$ uniformly on $\operatorname{supp}\mu$. Note that, since $X$ is compact, we have that $r_{k}(x)\leq\lvert X\rvert$, whenever $r_{k}(x)$ is defined.
@@
The conditions on the measure imply that, when defined implicitly, $r_{n}$ is well defined for each $n\in\mathbb{N}$ and that $r_{n}$ Lipschitz continuous on $\operatorname{supp}\mu$ with $\lvert r_{n}\rvert_{\textup{Lip}}=1$. In the case that $M_{n}\to 0$, then $r_{n}\to 0$ uniformly on $\operatorname{supp}\mu$. (The proofs of these facts are straightforward [38, Lemmas 3.1, 3.2].) When $r_{n}\to 0$ uniformly on $\operatorname{supp}\mu$, we may therefore without loss of generality assume that $r_{n}\leq\rho_{0}$ $\mu$-a.e. for all $n\in\mathbb{N}$.
@@
Conditions M.1 and M.2 are standard assumptions; they are assumed, either implicitly or explicitly, in the aforementioned works establishing RSBCs [16, 22, 29, 35, 36, 38]. If $\mu$ is absolutely continuous with density $h\in L^{q}(m)$ for some $q>1$, then both conditions hold. Furthermore, if $X$ is a Riemannian manifold with dimension $\dim X$ and $\mu$ satisfies Condition M.1 for $s>\dim X-1$, then Condition M.2 holds for $\alpha_{0}=\dim X+1-s$ (see the proof of Theorem 1.1 in [38]).
@@ 2.2. Distributional Limit Laws for Recurrence
  2. (2)
    The condition (2.3) that the ratio of variances converge in distribution can be verified if, for example, the system satisfies a short returns estimate of the following form: there exists $\delta_{0}>0$ such that, for all $k,j\geq 1$,
@@
    The condition (2.3) will be proven in Lemma 3.4 under a slightly weaker short returns estimate. The short returns estimate will be discussed further in Section 9.
Theorem 2.1 can be applied to interval maps equipped with a Gibbs measure and Axiom A systems equipped with SRB measures satisfying (M.1) for $s>\dim X-1$.
We also establish a distributional limit for the explicit radii problem. Recall that $\hat{E}_{n}=\{x\in X:T^{n}x\in B(x,r_{n})\}$ and $\hat{M}_{n}=\int\mu(B(x,r_{n}))\mathop{}\!d\mu(x)$. Let $\hat{\sigma}_{n}^{2}=\operatorname{Var}(\sum_{k=1}^{n}\mathbf{1}_{\hat{E}_{k}})$ and $\hat{s}_{n}^{2}=\operatorname{Var}(\sum_{k=1}^{n}\mathbf{1}_{B(y,r_{k})}\circ T^{k})$.
@@
Suppose that $(X,T,\mu,d)$ is an m.m.p.s. satisfying Condition D.3 and that $\mathop{}\!d\mu=h\mathop{}\!dm$ with density $h\in L^{q}(m)$ for some $q>2$ such that $h\geq c>0$ for some constant $c>0$. Assume that
@@
for all sufficiently large $k$ and all $l=1,\ldots,(\log k)^{p_{0}}$. If $(r_{n})$ satisfies Condition $\mathcal{S}_{\gamma,\upsilon}$ for some $\gamma\in(0,\dim X)$ and $\upsilon>2(\delta_{0}\dim X)^{-1}$, then the characteristic function of
@@
  1. (1)
    Although the result holds for $h\in L^{q}(m)$ for some $q>2$ given that (2.6) holds, we can only verify (2.6) when $h\in L^{\infty}(m)$. See §9 for further details.
  2. (2)
    Since $h$ is bounded away from 0, the limiting distribution has probability density function $f\colon\mathbb{R}\to(0,\infty)$ given by
    $f(t)=\frac{1}{\sqrt{2\pi}}\int_{X}\frac{\sqrt{\mu(h)}}{\sqrt{h(y)}}\exp\Bigl(-\frac{\mu(h)}{2h(y)}t^{2}\Bigr)\mathop{}\!d\mu(y)$
@@
  3. (3)
    Comparing with the result of Holland and Todd [22], namely (1.4), we see that both our probability density functions involve averaging over the ‘expected’ distributions. More precisely, the density in (1.4) is an average of Poisson distributions with variance $\tau h(x)$. In our case, our density is an average of Gaussian distributions with variance $h(x)/\mu(h)$.
@@ 2.3. Almost Sure Invariance Principles for Non-invertible Systems
As mentioned in the introduction, Haydn, Nicol, Vaienti and Zhang [14] prove an ASIP (and therefore a CLT) for the shrinking target problem when the transfer operator of the system contracts in the $\mathrm{BV}$ norm, i.e., under Condition D.1 with $\|\cdot\|_{\alpha}$ replaced with $\|\cdot\|_{\textup{BV}}$. Thus, their result can be used to establish the conclusion of Theorem 2.1 for such systems; however, in order to prove the conclusion for a larger class of systems – including uniformly hyperbolic systems – one needs to weaken their assumptions. In [14, Theorem 6.1], the authors assume, for their sequence of observables $(\varphi_{n})$, that $\sup_{n\geq 1}\|\varphi_{n}\|_{\alpha}<\infty$. Therefore, they are unable to apply their result to the shrinking target problem. Theorem 2.3 relaxes this condition, which enables us to use an approximation argument and establish an ASIP for the shrinking target problem. Additionally, we improve [14, Theorem 6.1] by weakening their condition that $\sigma_{n}^{2}\geq n^{\delta}$ for some $\delta>\frac{\sqrt{17}-1}{4}\approx 0.781$ to $\sigma_{n}^{2}\geq n^{\delta}$ for any $\delta>0$. (We will see that $\sigma_{n}^{2}\gtrsim n^{1-\gamma}$.)